Given y- y - 4y-6y=0 (1) , identify Differential Operator Lof (1). o L = D3 - D2 - 4D - 6 O None of them o L = (D - 3)(D
Answer #1


by Givern, y!_y_44-6y=0 (1) put y = by By - sty-you-Ey=0 D 4076 3-0 6.) 6 orod (0-3) (0+20+2) = ²5% 40-6=0 so, Differenti